#1 Cash Payback Period, Net Present Value Method, and Analysis for a Service Company

Social Circle Publications Inc. is considering two new magazine products. The estimated net cash flows from each product are as follows:

Year Sound Cellar Pro Gamer
1 $ 65,000 $ 70,000
2    60,000    55,000
3    25,000    35,000
4    25,000    30,000
5    45,000    30,000
Total $220,000 $220,000

Each product requires an investment of $125,000. A rate of 10% has been selected for the net present value analysis.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1a. Compute the cash payback period for each product.

Cash Payback Period
Sound Cellar
Pro Gamer

1b. Compute the net present value. Use the present value of $1 table above. If required, round to the nearest dollar.

Sound Cellar Pro Gamer
Present value of net cash flow total $ $
Amount to be invested $ $
Net present value $ $

2. Because of the timing of the receipt of the net cash flows, the   magazine expansion offers a higher  .

#2 Net Present Value Method, Internal Rate of Return Method, and Analysis for a Service Company

The management of Style Networks Inc. is considering two TV show projects. The estimated net cash flows from each project are as follows:

Year After Hours Sun Fun
1 $320,000 $290,000
2   320,000   290,000
3   320,000   290,000
4   320,000   290,000

After Hours requires an investment of $913,600, while Sun Fun requires an investment of $880,730. No residual value is expected from either project.

Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.352 2.991
6 4.917 4.355 4.111 3.784 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

Required:

1a. Compute the net present value for each project. Use a rate of 10% and the present value of an annuity of $1 in the above table. If required, round to the nearest dollar.

After Hours Sun Fun
Present value of annual net cash flows $ $
Amount to be invested $ $
Net present value $ $

1b. Compute a present value index for each project. If required, round your answers to two decimal places.

Present Value Index
After Hours
Sun Fun

2. Determine the internal rate of return for each project by (a) computing a present value factor for an annuity of $1 and (b) using the present value of an annuity of $1 table above. If required, round your present value factor answers to three decimal places and internal rate of return to the nearest percent.

After Hours Sun Fun
Present value factor for an annuity of $1
Internal rate of return   %   %

3. The net present value, present value index, and internal rate of return all indicate that the   TV show is a better financial opportunity compared to the   TV show, although both investments meet the minimum return criterion of 10%.
Answer #1

Social Circle Publications Inc.

1a.

Payback period:
Sound Cellar 2 year
Pro Gamar 2 year

Working:

Sound Pro Payback Amount Payback period
Year Celler Gammer Celler Gammer Celler Gammer
1 $       65,000.00 $      70,000.00 $ 65,000.00 $ 70,000.00 1 1
2 $       60,000.00 $      55,000.00 $ 60,000.00 $ 55,000.00 1 1
3 $       25,000.00 $      35,000.00
4 $       25,000.00 $      30,000.00
5 $       45,000.00 $      30,000.00
Total $     220,000.00 $     220,000.00 $125,000.00 $125,000.00 2 2

1b.

Sound Pro
Celler Gammer
Present Value of Net Cash flow 172440 174465
Less: Amount Invested 125000 125000
Net Present Value 47440 49465

Working:

Sound Pro PV Factor Present Value
Year Celler Gammer at 10% Celler Gammer
1 65000 70000 0.909 59085 63630
2 60000 55000 0.826 49560 45430
3 25000 35000 0.751 18775 26285
4 25000 30000 0.683 17075 20490
5 45000 30000 0.621 27945 18630
Total 220000 220000 172440 174465

2.

Because of the timing of the receipt of the net cash flows, the Pro Gamar magazine expenses offer a higher Net Present Value.

Style Network Inc.

1a.

1a. NPV = sum of all present values (PV). PV = amount of cash flow*present value of annuity at 10% for the applicable year.
NPV of after hours = (present value of cash flows from years 1 to 4) - initial investment.
present value of cash flows from years 1 to 4 = annual cash flows of 320,000*PV of annuty at 10% and 4 years = 320,000*3.17 = $1,014,400.
Thus NPV = $1,014,400 - 913,600 = $100,800
Sun Fun: present value of cash flows from years 1 to 4 = 290,000*3.17 = $919,300. Thus NPV = 919,300 - 880,730 = 38,750
After Hours Sun Fun
Present value of annual net cash flows 1,014,400.00 919,300.00
Less amount to be invested -913,600.00 -880,730.00
Net present value 100,800.00 38,570.00

1b.

1b. Present value Index = Present value of future cash flows/Initial investment
After hours = 1,014,400/913,600 = 1.11
Sun Fun = 919,300/880,730 = 1.04
Present Value Index
After Hours 1.11
Sun Fun 1.04

2.

2. IRR (internal rate of return) is that rate which makes the NPV as nil.
Now this is found using the trial and error method.
For After Hours the IRR is around 15%. From the provided table the PV factor of annuity = 2.855
Similarly for Sun Fun the IRR is 12% and PV factor of annuity = 3.037
After Hours Sun Fun
Present value factor for an annuity of $1 2.855 3.037

Internal ra

te of return

15 12

3.

The net present value, present value index, and internal rate of return all indicate that the After Hours TV

show is a better financial opportunity compared to the Sun Fun TV show.